# Context-Free Recognition with Transformers

### v1

@@ B.2 Unambiguous CFL recognition on transformers
Solving reachability queries:
Because the given grammar is unambiguous, for each node $v$, the subgraph induced by nodes reachable from $v$ becomes a tree rooted at $v$ (Fact˜4.1). Reachability queries over binary trees now reduce to evaluating the Boolean formula associated with the binary tree. Leaf nodes associated with realizable items are assigned $\mathrm{T}$}. A non-leaf node has a path to such a leaf if evaluating the induced Boolean expression where non-leaf compute $\lor$ over their children yields $\mathrm{T}$}. We can therefore invoke Lem.˜4.1 to evaluate this Boolean formula.
Recognition step:
The eos token can attend to the padding token for node associated with ${{{{\mathrm{S}}}}},1,{{n}}]$ and check whether it is realizable, i.e., store $1$ in its residual stream. ∎
@@ Appendix C Experimental setup
We used butoi-languageclasses’s length-constrained sampling algorithm for CFLs to generate datasets. For $\mathrm{D}\!\left(1\right)$}}, $\mathrm{D}\!\left(2\right)$}}, Palindrome and Marked Palindrome, negative samples were either sampled at random from $\Sigma$$^{*}$}} or were perturbations from positive strings. For BFVP, the negative strings were sampled Boolean formulas that evaluate to $\mathrm{F}$} as we preferred to focus on a transformer’s ability to correctly evaluate a Boolean formula rather than determining if the formula is well-formed. The ability to process hierarchically nested structures is already captured by the language $\mathrm{D}\!\left({{k}}\right)$}}. The training set consists of 1 million samples with string length at most 40. The test set has 2000 samples with string length at most 80. Testing the model on strings longer than those seen in training enabled the evaluation of its ability to generalize out-of-distribution.

### v2

@@ B.2 Unambiguous CFL Recognition on Transformers
Solving reachability queries. Because the given grammar is unambiguous, for each node $v$, the subgraph induced by nodes reachable from $v$ becomes a tree rooted at $v$ (Fact˜4.1). Reachability queries over binary trees now reduce to evaluating the Boolean formula associated with the binary tree. Leaf nodes associated with realizable items are assigned $\mathrm{T}$}. A non-leaf node has a path to such a leaf if evaluating the induced Boolean expression where non-leaf compute $\lor$ over their children yields $\mathrm{T}$}. We can therefore invoke Lem.˜4.1 to evaluate this Boolean formula.
Recognition step. The eos token can attend to the padding token for node associated with ${{{{\mathrm{S}}}}},1,{{n}}]$ and check whether it is realizable, i.e., store $1$ in its residual stream. ∎
@@ Appendix C Experimental Setup
We used butoi-languageclasses’s length-constrained sampling algorithm for CFLs to generate datasets. To sample strings from a given grammar $\mathcal{G}$}}, we consider the probabilistic version of $\mathcal{G}$}} which induces a probability distribution over ${\mathbb{L}}({{\mathcal{G}}})$, which enables length-constrained sampling. Importantly, their procedure first samples a desired string length n{{n}}, and then performs sampling over the distribution of all strings of length n{{n}}. Negative strings are either sampled at random from $\Sigma$$^{*}$}} or are perturbations from positive strings by applying random edits to them, the number of which is randomly sampled from a geometric distribution that favors small values (butoi2025trainingneuralnetworksrecognizers).
We therefore follow this procedure to sample positive and negative strings from handpicked context-free grammars except for BFVP. For BFVP, the negative strings were sampled Boolean formulas that evaluate to $\mathrm{F}$}, such that the transformer is trained to correctly evaluate a Boolean formula rather than checking if an input is well-formed. We argue that the ability to process hierarchically nested structures, such as nested subformulas in BFVP, is already captured by the grammar $\mathrm{D}\!\left({{k}}\right)$}}.
The training set consists of 1 million samples with string length at most 40. The test set has 2000 samples with string length at most 80. Testing the model on strings longer than those seen in training enabled the evaluation of its ability to generalize out-of-distribution.